Part I: You, the accountant, are analyzing Nolans Corporation. Nolans corp. has expanded its production facilities by 200% since 2016. The income statements for the last three year are below:

Nolan Security Corporation

Statement of Financial Position

At December 31, 2020

Assets

Liabilities

Current

Current

Cash

$100

Accounts Payable

$300

Accounts Receivable

200

Wages Payable

50

Merchandise Inventory

500

Dividends Payable

50

Prepared Expenses

50

400

850

Non-current

Non-current

   Property, plant & equipment (net)

1,000

Loan Payable

800

1,200

Shareholders’ Equity

Common Shares

500

Retained Earnings

150

Total Shareholder’s Equity

650

Total Assets

$1,850

Total Liability and Equity

$1,850

Requirements: Show all work.

1. Calculate current ratio, acid-test ratio, and debt to shareholder’s equity ratio.

2. Discuss the meaning of each ratio result. (What do these ratios mean for Nolans Security?)

Part III: The following balance sheet (statement of financial position) is presented for LevelUp Corporation.

LevelUp Corporation

Statement of Financial Position

At December 31, 2020

Assets

Liabilities

Current

Current

Cash

$60

Accounts Payable

$100

Accounts Receivable

140

Loan Payable

20

Merchandise Inventory

250

Notes Payable

60

Prepared Expenses

10

180

460

Non-current

Non-current

   Property, plant & equipment (net)

330

Loan Payable

140

320

Shareholders’ Equity

Preferred shares, 10% (8 shares)

120

Common shares (50 shares)

250

Retained earnings

100

470

Total Assets

$790

Total Liability and Equity

$790

LevelUp Corporation

Income Statement

For the Year Ending December 31, 2020

Net Sales (all on credit)

$800

Cost of Goods Sold

600

Gross Profit

200

Selling and Administration Expenses

100

Income from Operations

100

Interest Expense

20

Income before Income Taxes

80

Income Taxes

30

Net Income

$50

Additional information from December 31, 2019 statement of financial position:

Accounts receivable                            $180

Merchandise inventory                        200

Property, Plant and Equipment (net) 250

Retained earnings                                   80

Preferred shares                                  120

Common Shares                                  250   

Requirements:

1. Compute the following ratios, showing all work.

Current ratio

Acid-test ratio

Accounts receivable collection period

Number of days of sales in inventory

Debt to shareholders’ equity ratio

Return on shareholder’s equity

2. What do these ratios tell you about the LevelUp Corporation?
Answer #1

Part-1

Current ratio=current asset/current liabilities

Current asset=Cash+account receivable+merchandise inventory+prepaid expenses

=100+200+500+50=$850

Current liabilities= Account payable+ wages payable+ dividend payable

=300+50+50=$400

Therefore,

Current ratio= $850/$400=2.125:1

Acid test ratio= (current asset-inventory-prepaid expenses)/current liabilities

=($850-$500-$50)/400= 0.75:1

Debt to shareholders equity ratio= Debt/shareholders equity

= Loan payable/shareholders equity

= 800/650=1.23:1

Current ratio is adequate in this case as ideal current ratio is 2:1

Acid test ratio is not adequate as ideal ratio is 1:1, asset are less than liabilities

Debt equity ratio is adequate as ideal ratio lies between 1 to 1.5.

Part-III

Current ratio=current asset/current liabilities

Current asset=Cash+account receivable+merchandise inventory+prepaid expenses

=60+140+250+10=$460

Current liabilities= Account payable+ Loan payable + notes payable

=100+20+60=$180

Therefore,

Current ratio= $460/$180=2.55:1

Acid test ratio= (current asset-inventory-prepaid expenses)/current liabilities

=($460-$250-$10)/180= 1.11:1

Account receivable collection period= (Account receivable/sales)*365

= ($140/$800)*365= 63.875 days or 64 days(rounded off)

Number of days of sales in inventory= ( inventory/COGS)*365

=($250/$600)*365= 152 days approx.

Debt to shareholders equity= Debt/shareholders equity

=$140/$470= 0.30:1

Return on shareholders equity= net income/ shareholders equity*100

= ($50/$470)*100=10.64%

Current ratio adequate as it is above ideal ratio of 2:1

Acid test ratio is adequate as it is above ideal ratio of 1:1

Account receivable collection period is adequate as ideal period is of 45 days

Debt to shareholders equity is not appropriate as it is below ideal ratio.

No. Of days of sale in inventory is 152 days which shows that this much time company takes to convert inventory into cash.

Return on equity is not adequate as ideal percentage should be in between 15%-20%.